OpenStudy (anonymous):

is the statement correct,,,,,,,,,,,,plse help me Construct a triangle abc in which bc = 5 cm , angle c =60 and ac-ab = 1.5 cm

OpenStudy (raden):

just say true or false ? i didnt understand with the question --"

OpenStudy (anonymous):

ac-ab = 1.5 cm is written correct or not or it should be ab-ac=1.5 cm

OpenStudy (raden):

hmmm... i caught it, we need prove is this triangle or not according the values given above. let AB = x AC = x + 1.5 cos60 = [5^2 + (x+1.5)^2 - x^2]/(2*5*(x+1.5)) 1/2 = [25+3x+2.25]/(10x+15) 10x+15 = 55+6x 4x = 40 x = 10 so, AB = 10, AC=11.5, and BC=5 whatever, it is a triangle, because : 10+11.5 > 5 10+5 > 11.5 5+11.5 > 10 (MAYBE)

OpenStudy (anonymous):

k thank u

OpenStudy (raden):

you're welcome